OpenStudy (anonymous):

Please explain this simply for me! Medical x-rays, with a wavelength of about 10^(-10) m, can penetrate completely through your skin. a. Ultraviolet rays, which cause sun burn by penetrating only the top level of skin, have a wavelength about 1,000 times the wavelength of an x-ray. Find the wavelength of ultraviolet rays. Show your work.

OpenStudy (amistre64):

medwave*1000 id assume

OpenStudy (lgbasallote):

10^-10 x 1000 100 = 10^3 so 10^-10 x 10^3 add exponents... 10^-7 m am i right?

OpenStudy (amistre64):

looks right to me

OpenStudy (anonymous):

Thanks! There's a second part - b. The wavelengths of visible light are between 3.8 x 10^(-7) m and 7.6 x 10^(-7) m. Are these wavelengths longer or shorter than the ultraviolet rays? Explain.

OpenStudy (lgbasallote):

ultraviolet rays is 0.000010 m 0.00000038<visible light<0.00000076 which do you think is longer/shorter?

OpenStudy (anonymous):

Ultraviolet rays?

OpenStudy (anonymous):

*are longer.

OpenStudy (lgbasallote):

yeah...i think so too...but you're the med student here :D

OpenStudy (anonymous):

XD I'm not a med student, I'm just in Algebra 1...

OpenStudy (lgbasallote):

of all the examples why this -____- your teacher is good....at torture >:))

OpenStudy (anonymous):

I know, I know. and I hate Algebra on top of this all, it's confusing. The next question confuses me even more! DX

OpenStudy (lgbasallote):

try me ;D

OpenStudy (anonymous):

I'll post it...

OpenStudy (jamesj):

Wait a second. You just calculated that UV light has wavelengths of \( 10^{-7} \) meters. That is: \[ 1 \times 10^{-7} \ m \] Then you're told that visible light has wavelengths such as \[ 3.8 \times 10^{-7} m \] and \[ 7.6 \times 10^{-7} \ m \] Clearly therefore, because 3.8 > 1 and 7.6 > 1, visible light has a longer wavelength.

OpenStudy (anonymous):

But the lengths that are being compared are ultraviolet and medical x-ray.

OpenStudy (lgbasallote):

james has a point...

OpenStudy (jamesj):

Yes. I just showed you why visible light has a longer wave length. Do you agree with that? Then because \[ 10^{-10} < 10^{-7} \] it must be that x-rays also have a shorter wave length than visible light.

OpenStudy (anonymous):

Ohhh. I think I see...